Imagine that you are a Labor Relations Manager employed with a private, for-profit company preparing for an upcoming labor contract negotiation. As a part of their integrative bargaining strategy, senior management has considered opening the contract talks with a generous offer of a 3% across the board wage increase. This is expected to be accepted by the union negotiating team in exchange for concessions for flexible hiring. Senior management has instructed you to calculate the impact of this wage increase and report back.

Your current union workforce data is as follows:

97 employees at $11.25/hour             Average number of work hours per year = 2000

69 employees at $13.50/hour             Average number of paid vacation hours per year = 80

34 employees at $15.75/hour             Average cost of benefits per person = $4.73/hour or 35%

Using the data above calculate the following:

1a. What is the current annual cost of employee base wages?

1b. What will the annual cost of employee base wages be, if proposed increase is accepted?

1c. How much additional wage cost will be added by this proposed increase?

2a. What is the current total annual cost of employee benefits?

2b. What will the total annual cost of benefits be, if proposed increase is accepted?

2c. How much additional benefit cost will be added by this wage increase?

3. What is the total cost of this proposed wage increase?
Answer #1

The current annual base wage cost is considered considering work hr and vacation hrs.

Wage Benefits Total cost/hr Employee Work hr+ Vacation Total base Wages cost Total wage cost @3% Total cost including benefit' Total cost including benefit' with increase
11.25 4.73 15.98 97 2080 2269800 2337894 3224124.8 3320848.544
13.5 4.73 18.23 69 2080 1937520 1995645.6 2616369.6 2694860.688
15.75 4.73 20.48 34 2080 1113840 1147255.2 1448345.6 1491795.968
5321160 5480794.8 7288840 7507505.2

1.a) Total base wage cost= Σ of 97, 69, 34(wage* number of employee*(average hr+ vacation hrs)= 5321160

b) Total base wage at increased cost=5480794

c) Increase in wage- b-a= 159634

2.a) Current cost due to benefit= Annual cost with benefit- annual cost without benefit=7288840-5321160= 1967680

2. b) Cost of benefit if proposed increase is accepted= 7507505-5480794= 2026710

2.c) Additional benefit added=2026710-1967680=59030

3) Difference between the overall change= 2026710